OpenStudy (anonymous):

log2(x-1)+log2(x+1)=3 How do I work this problem out? Answer should be 3.

OpenStudy (anonymous):

Is the 2 after the log the base or is it 2 multiplied by the quantity?

OpenStudy (anonymous):

log base of 2, not mulitply

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

when adding logs you can bring them together by multiplying there arguments. \[\log_{2}[(x-1)*(x+1)]=3\] \[(x-1)(x+1)=2^3\] \[x^2-1=8\] \[x=\pm3\]

OpenStudy (anonymous):

thank you i appreciate your time. i will practice these questions after i get rest :P

OpenStudy (anonymous):

yea no problem. =)